# Optimized effective potential forces with the plane-wave and pseudopotential method

### v1

@@ Acknowledgements.
- Makmal et al. [2009a] A. Makmal, S. Kümmel, and L. Kronik, Journal of Chemical Theory and Computation 5, 1731 (2009a).
- Hellgren and von Barth [2008] M. Hellgren and U. von Barth, Phys. Rev. B 78, 115107 (2008).
- Hellgren and von Barth [2009] M. Hellgren and U. von Barth, The Journal of Chemical Physics 131, 044110 (2009).
- Hellgren and von Barth [2010] M. Hellgren and U. von Barth, The Journal of Chemical Physics 132, 044101 (2010).
- Hellgren and Gross [2013] M. Hellgren and E. K. U. Gross, Phys. Rev. A 88, 052507 (2013).
- Yang and Wu [2002] W. Yang and Q. Wu, Phys. Rev. Lett. 89, 143002 (2002).
- Kümmel and Perdew [2003a] S. Kümmel and J. P. Perdew, Phys. Rev. Lett. 90, 043004 (2003a).
- Kümmel and Perdew [2003b] S. Kümmel and J. P. Perdew, Phys. Rev. B 68, 035103 (2003b).
- Nguyen et al. [2014] N. L. Nguyen, N. Colonna, and S. de Gironcoli, Phys. Rev. B 90, 045138 (2014).
- Sharma et al. [2005] S. Sharma, J. K. Dewhurst, and C. Ambrosch-Draxl, Phys. Rev. Lett. 95, 136402 (2005).
@@
- Makmal et al. [2009b] A. Makmal, R. Armiento, E. Engel, L. Kronik, and S. Kümmel, Phys. Rev. B 80, 161204(R) (2009b).
- Betzinger et al. [2011] M. Betzinger, C. Friedrich, S. Blügel, and A. Görling, Phys. Rev. B 83, 045105 (2011).
- Kümmel and Kronik [2008] S. Kümmel and L. Kronik, Rev. Mod. Phys. 80, 3 (2008).
- Trushin and Görling [2018] E. Trushin and A. Görling, The European Physical Journal B 91, 149 (2018).
- Trushin et al. [2019] E. Trushin, L. Fromm, and A. Görling, Phys. Rev. B 100, 075205 (2019).
- Rinke et al. [2005] P. Rinke, A. Qteish, J. Neugebauer, C. Freysoldt, and M. Scheffler, New Journal of Physics 7, 126 (2005).
- Qteish et al. [2006] A. Qteish, P. Rinke, M. Scheffler, and J. Neugebauer, Phys. Rev. B 74, 245208 (2006).
@@
- Wu et al. [2005] Q. Wu, A. J. Cohen, and W. Yang, The Journal of Chemical Physics 123, 134111 (2005).
- Thierbach and Görling [2020] A. Thierbach and A. Görling, The Journal of Chemical Physics 152 (2020).
- Thierbach and Görling [2020] A. Thierbach and A. Görling, The Journal of Chemical Physics 153, 134113 (2020).
- Krieger et al. [1992b] J. B. Krieger, Y. Li, and G. J. Iafrate, Phys. Rev. A 45, 101 (1992b).
@@
- Sham and Schlüter [1983] L. J. Sham and M. Schlüter, Phys. Rev. Lett. 51, 1888 (1983).
- Casida [1995] M. E. Casida, Phys. Rev. A 51, 2005 (1995).
@@
- Giannozzi et al. [2017] P. Giannozzi, O. Andreussi, T. Brumme, O. Bunau, M. B. Nardelli, M. Calandra, R. Car, C. Cavazzoni, D. Ceresoli, M. Cococcioni, N. Colonna, I. Carnimeo, A. D. Corso, S. de Gironcoli, P. Delugas, R. A. DiStasio, A. Ferretti, A. Floris, G. Fratesi, G. Fugallo, R. Gebauer, U. Gerstmann, F. Giustino, T. Gorni, J. Jia, M. Kawamura, H.-Y. Ko, A. Kokalj, E. Kucukbenli, M. Lazzeri, M. Marsili, N. Marzari, F. Mauri, N. L. Nguyen, H.-V. Nguyen, A. O. de-la Roza, L. Paulatto, S. Poncé, D. Rocca, R. Sabatini, B. Santra, M. Schlipf, A. P. Seitsonen, A. Smogunov, I. Timrov, T. Thonhauser, P. Umari, N. Vast, X. Wu, and S. Baroni, Journal of Physics: Condensed Matter 29, 465901 (2017).
@@
- Baroni et al. [1987] S. Baroni, P. Giannozzi, and A. Testa, Phys. Rev. Lett. 58, 1861 (1987).
- Hellgren and Baguet [2021] M. Hellgren and L. Baguet, Phys. Rev. Res. 3, 033263 (2021).
- Baroni et al. [2001] S. Baroni, S. de Gironcoli, A. Dal Corso, and P. Giannozzi, Rev. Mod. Phys. 73, 515 (2001).
@@
- Pick et al. [1970] R. M. Pick, M. H. Cohen, and R. M. Martin, Phys. Rev. B 1, 910 (1970).
- Hellman [1937] H. Hellman, Franz Deuticke, Leipzig 285 (1937).
- Feynman [1939] R. P. Feynman, Phys. Rev. 56, 340 (1939).
@@
- Hamann et al. [1979] D. R. Hamann, M. Schlüter, and C. Chiang, Phys. Rev. Lett. 43, 1494 (1979).
@@
- Kleinman and Bylander [1982] L. Kleinman and D. M. Bylander, Phys. Rev. Lett. 48, 1425 (1982).
- Blöchl [1990] P. E. Blöchl, Phys. Rev. B 41, 5414 (1990).
- Vanderbilt [1990] D. Vanderbilt, Phys. Rev. B 41, 7892 (1990).
- Nguyen and de Gironcoli [2009] H.-V. Nguyen and S. de Gironcoli, Phys. Rev. B 79, 205114 (2009).
- Broyden [1970] C. G. Broyden, IMA Journal of Applied Mathematics 6, 76 (1970).
- Fletcher [1970] R. Fletcher, The Computer Journal 13, 317 (1970).
- Goldfarb [1970] D. Goldfarb, Math. Comp. 24, 23 (1970).
- Shanno [1970] D. F. Shanno, Math. Comp. 24, 647 (1970).
- Hamann [2013] D. R. Hamann, Phys. Rev. B 88, 085117 (2013).
- Thomas et al. [1993] J. R. Thomas, B. J. DeLeeuw, G. Vacek, T. D. Crawford, Y. Yamaguchi, and I. Schaefer, Henry F., The Journal of Chemical Physics 99, 403 (1993).
- Clabo et al. [1988] D. Clabo, W. D. Allen, R. B. Remington, Y. Yamaguchi, and H. F. Schaefer, Chemical Physics 123, 187 (1988).
- Togo et al. [2023] A. Togo, L. Chaput, T. Tadano, and I. Tanaka, Journal of Physics: Condensed Matter 35, 353001 (2023).
- Togo [2023] A. Togo, Journal of the Physical Society of Japan 92, 012001 (2023).
- Levien et al. [1980] L. Levien, C. T. Prewitt, and D. J. Weidner, American Mineralogist 65, 920 (1980).
- Mounet and Marzari [2005] N. Mounet and N. Marzari, Phys. Rev. B 71, 205214 (2005).
- He et al. [2014] L. He, F. Liu, G. Hautier, M. J. T. Oliveira, M. A. L. Marques, F. D. Vila, J. J. Rehr, G.-M. Rignanese, and A. Zhou, Phys. Rev. B 89, 064305 (2014).
- Gonze and Lee [1997] X. Gonze and C. Lee, Phys. Rev. B 55, 10355 (1997).

### v2

@@ Acknowledgements.
- Makmal et al. [2009a] A. Makmal, S. Kümmel, and L. Kronik, Fully Numerical All-Electron Solutions of the Optimized Effective Potential Equation for Diatomic Molecules, Journal of Chemical Theory and Computation 5, 1731-1740 (2009a).
- Hellgren and von Barth [2008] M. Hellgren and U. von Barth, Linear density response function within the time-dependent exact-exchange approximation, Phys. Rev. B 78, 115107 (2008).
- Hellgren and von Barth [2009] M. Hellgren and U. von Barth, Exact-exchange kernel of time-dependent density functional theory: Frequency dependence and photoabsorption spectra of atoms, The Journal of Chemical Physics 131, 044110 (2009).
- Hellgren and von Barth [2010] M. Hellgren and U. von Barth, Correlation energy functional and potential from time-dependent exact-exchange theory, The Journal of Chemical Physics 132, 044101 (2010).
- Hellgren and Gross [2013] M. Hellgren and E. K. U. Gross, Discontinuous functional for linear-response time-dependent density-functional theory: The exact-exchange kernel and approximate forms, Phys. Rev. A 88, 052507 (2013).
- Yang and Wu [2002] W. Yang and Q. Wu, Direct Method for Optimized Effective Potentials in Density-Functional Theory, Phys. Rev. Lett. 89, 143002 (2002).
- Kümmel and Perdew [2003a] S. Kümmel and J. P. Perdew, Simple Iterative Construction of the Optimized Effective Potential for Orbital Functionals, Including Exact Exchange, Phys. Rev. Lett. 90, 043004 (2003a).
- Kümmel and Perdew [2003b] S. Kümmel and J. P. Perdew, Optimized effective potential made simple: Orbital functionals, orbital shifts, and the exact Kohn-Sham exchange potential, Phys. Rev. B 68, 035103 (2003b).
- Nguyen et al. [2014] N. L. Nguyen, N. Colonna, and S. de Gironcoli, Ab initio self-consistent total-energy calculations within the EXX/RPA formalism, Phys. Rev. B 90, 045138 (2014).
- Engel et al. [2001] E. Engel, A. Höck, R. N. Schmid, R. M. Dreizler, and N. Chetty, Role of the core-valence interaction for pseudopotential calculations with exact exchange, Phys. Rev. B 64, 125111 (2001).
- Sharma et al. [2005] S. Sharma, J. K. Dewhurst, and C. Ambrosch-Draxl, All-Electron Exact Exchange Treatment of Semiconductors: Effect of Core-Valence Interaction on Band-Gap and -Band Position, Phys. Rev. Lett. 95, 136402 (2005).
@@
- Makmal et al. [2009b] A. Makmal, R. Armiento, E. Engel, L. Kronik, and S. Kümmel, Examining the role of pseudopotentials in exact-exchange-based Kohn-Sham gaps, Phys. Rev. B 80, 161204(R) (2009b).
- Betzinger et al. [2011] M. Betzinger, C. Friedrich, S. Blügel, and A. Görling, Local exact exchange potentials within the all-electron FLAPW method and a comparison with pseudopotential results, Phys. Rev. B 83, 045105 (2011).
- Kümmel and Kronik [2008] S. Kümmel and L. Kronik, Orbital-dependent density functionals: Theory and applications, Rev. Mod. Phys. 80, 3–60 (2008).
- Trushin and Görling [2018] E. Trushin and A. Görling, Assessment of quality and reliability of band structures from exact-exchange-only Kohn–Sham, hybrid, and GW methods, The European Physical Journal B 91, 149 (2018).
- Trushin et al. [2019] E. Trushin, L. Fromm, and A. Görling, Assessment of the exact-exchange-only Kohn-Sham method for the calculation of band structures for transition metal oxide and metal halide perovskites, Phys. Rev. B 100, 075205 (2019).
- Rinke et al. [2005] P. Rinke, A. Qteish, J. Neugebauer, C. Freysoldt, and M. Scheffler, Combining calculations with exact-exchange density-functional theory: an analysis of valence-band photoemission for compound semiconductors, New J. Phys. 7, 126 (2005).
- Qteish et al. [2006] A. Qteish, P. Rinke, M. Scheffler, and J. Neugebauer, Exact-exchange-based quasiparticle energy calculations for the band gap, effective masses, and deformation potentials of ScN, Phys. Rev. B 74, 245208 (2006).
@@
- Wu et al. [2005] Q. Wu, A. J. Cohen, and W. Yang, Analytic energy gradients of the optimized effective potential method, The Journal of Chemical Physics 123, 134111 (2005).
- Thierbach and Görling [2020] A. Thierbach and A. Görling, Analytic energy gradients for the exact exchange Kohn–Sham method, The Journal of Chemical Physics 152, 114113 (2020).
- Thierbach and Görling [2020] A. Thierbach and A. Görling, Analytic energy gradients for the self-consistent direct random phase approximation, The Journal of Chemical Physics 153, 134113 (2020).
@@
- Sham and Schlüter [1983] L. J. Sham and M. Schlüter, Density-Functional Theory of the Energy Gap, Phys. Rev. Lett. 51, 1888–1891 (1983).
- Casida [1995] M. E. Casida, Generalization of the optimized-effective-potential model to include electron correlation: A variational derivation of the Sham-Schlüter equation for the exact exchange-correlation potential, Phys. Rev. A 51, 2005–2013 (1995).
@@
- Giannozzi et al. [2017] P. Giannozzi, O. Andreussi, T. Brumme, O. Bunau, M. B. Nardelli, M. Calandra, R. Car, C. Cavazzoni, D. Ceresoli, M. Cococcioni, N. Colonna, I. Carnimeo, A. D. Corso, S. de Gironcoli, P. Delugas, R. A. DiStasio, A. Ferretti, A. Floris, G. Fratesi, G. Fugallo, R. Gebauer, U. Gerstmann, F. Giustino, T. Gorni, J. Jia, M. Kawamura, H.-Y. Ko, A. Kokalj, E. Kucukbenli, M. Lazzeri, M. Marsili, N. Marzari, F. Mauri, N. L. Nguyen, H.-V. Nguyen, A. O. de-la Roza, L. Paulatto, S. Poncé, D. Rocca, R. Sabatini, B. Santra, M. Schlipf, A. P. Seitsonen, A. Smogunov, I. Timrov, T. Thonhauser, P. Umari, N. Vast, X. Wu, and S. Baroni, Advanced capabilities for materials modelling with Quantum ESPRESSO, Journal of Physics: Condensed Matter 29, 465901 (2017).
@@
- Baroni et al. [1987] S. Baroni, P. Giannozzi, and A. Testa, Green’s-function approach to linear response in solids, Phys. Rev. Lett. 58, 1861–1864 (1987).
- Hellgren and Baguet [2021] M. Hellgren and L. Baguet, Random phase approximation with exchange for an accurate description of crystalline polymorphism, Phys. Rev. Res. 3, 033263 (2021).
- Baroni et al. [2001] S. Baroni, S. de Gironcoli, A. Dal Corso, and P. Giannozzi, Phonons and related crystal properties from density-functional perturbation theory, Rev. Mod. Phys. 73, 515–562 (2001).
@@
- Pick et al. [1970] R. M. Pick, M. H. Cohen, and R. M. Martin, Microscopic Theory of Force Constants in the Adiabatic Approximation, Phys. Rev. B 1, 910–920 (1970).
- Hellman [1937] H. Hellman, Einführung in die Quantenchemie, Franz Deuticke, Leipzig 285 (1937).
- Feynman [1939] R. P. Feynman, Forces in Molecules, Phys. Rev. 56, 340–343 (1939).
@@
- Hamann et al. [1979] D. R. Hamann, M. Schlüter, and C. Chiang, Norm-Conserving Pseudopotentials, Phys. Rev. Lett. 43, 1494–1497 (1979).
@@
- Kleinman and Bylander [1982] L. Kleinman and D. M. Bylander, Efficacious Form for Model Pseudopotentials, Phys. Rev. Lett. 48, 1425–1428 (1982).
- Blöchl [1990] P. E. Blöchl, Generalized separable potentials for electronic-structure calculations, Phys. Rev. B 41, 5414–5416 (1990).
- Vanderbilt [1990] D. Vanderbilt, Soft self-consistent pseudopotentials in a generalized eigenvalue formalism, Phys. Rev. B 41, 7892–7895 (1990).
- Nguyen and de Gironcoli [2009] H.-V. Nguyen and S. de Gironcoli, Efficient calculation of exact exchange and RPA correlation energies in the adiabatic-connection fluctuation-dissipation theory, Phys. Rev. B 79, 205114 (2009).
- Broyden [1970] C. G. Broyden, The Convergence of a Class of Double-rank Minimization Algorithms 1. General Considerations, IMA Journal of Applied Mathematics 6, 76-90 (1970).
- Fletcher [1970] R. Fletcher, A new approach to variable metric algorithms, The Computer Journal 13, 317-322 (1970).
- Goldfarb [1970] D. Goldfarb, A family of variable-metric methods derived by variational means, Math. Comp. 24, 23-26 (1970).
- Shanno [1970] D. F. Shanno, Conditioning of quasi-Newton methods for function minimization, Math. Comp. 24, 647-656 (1970).
- Hamann [2013] D. R. Hamann, Optimized norm-conserving Vanderbilt pseudopotentials, Phys. Rev. B 88, 085117 (2013).
- Thomas et al. [1993] J. R. Thomas, B. J. DeLeeuw, G. Vacek, T. D. Crawford, Y. Yamaguchi, and H. F. Schaefer, The balance between theoretical method and basis set quality: A systematic study of equilibrium geometries, dipole moments, harmonic vibrational frequencies, and infrared intensities, The Journal of Chemical Physics 99, 403-416 (1993).
- Clabo et al. [1988] D. A. Clabo, W. D. Allen, R. B. Remington, Y. Yamaguchi, and H. F. Schaefer, A systematic study of molecular vibrational anharmonicity and vibration—rotation interaction by self-consistent-field higher-derivative methods. Asymmetric top molecules, Chemical Physics 123, 187-239 (1988).
- Togo et al. [2023] A. Togo, L. Chaput, T. Tadano, and I. Tanaka, Implementation strategies in phonopy and phono3py, Journal of Physics: Condensed Matter 35, 353001 (2023).
- Togo [2023] A. Togo, First-principles Phonon Calculations with Phonopy and Phono3py, Journal of the Physical Society of Japan 92, 012001 (2023).
- Levien et al. [1980] L. Levien, C. T. Prewitt, and D. J. Weidner, Structure and elastic properties of quartz at pressure, American Mineralogist 65, 920-930 (1980).
- Mounet and Marzari [2005] N. Mounet and N. Marzari, First-principles determination of the structural, vibrational and thermodynamic properties of diamond, graphite, and derivatives, Phys. Rev. B 71, 205214 (2005).
- He et al. [2014] L. He, F. Liu, G. Hautier, M. J. T. Oliveira, M. A. L. Marques, F. D. Vila, J. J. Rehr, G.-M. Rignanese, and A. Zhou, Accuracy of generalized gradient approximation functionals for density-functional perturbation theory calculations, Phys. Rev. B 89, 064305 (2014).
- Gonze and Lee [1997] X. Gonze and C. Lee, Dynamical matrices, Born effective charges, dielectric permittivity tensors, and interatomic force constants from density-functional perturbation theory, Phys. Rev. B 55, 10355–10368 (1997).